OpenStudy (anonymous):

What is the least common denominator for the fractions 5/6 and 3/8?

OpenStudy (tkhunny):

6 = 2*3 8 = 2*2*2 That's all you need.

OpenStudy (anonymous):

The answer is 24

OpenStudy (anonymous):

probobly bro...

OpenStudy (anonymous):

@Happie you need to explain .. COC For (6, 8) the least common multiple (LCM) is 24. Therefore, the least common denominator (LCD) is 24. Rewriting as equivalent fractions with the LCD: 20/24, 9/24

OpenStudy (anonymous):

so yea its 24 then?